### v2

# T-GVC: Trajectory-Guided Generative Video Coding at Ultra-Low Bitrates

Abstract
Recent advances in video generation techniques have given rise to an emerging paradigm of generative video coding, aiming to achieve semantically accurate reconstructions in Ultra-Low Bitrate (ULB) scenarios by leveraging strong generative priors. However, most existing methods are limited by domain specificity (e.g., facial or human videos) or an excessive dependence on high-level text guidance, which often fails to capture motion details and results in unrealistic reconstructions. To address these challenges, we propose a Trajectory-Guided Generative Video Coding framework (dubbed T-GVC). T-GVC employs a semantic-aware sparse motion sampling pipeline to effectively bridge low-level motion tracking with high-level semantic understanding by extracting pixel-wise motion as sparse trajectory points based on their semantic importance, not only significantly reducing the bitrate but also preserving critical temporal semantic information. In addition, by incorporating trajectory-aligned loss constraints into diffusion processes, we introduce a training-free latent space guidance mechanism to ensure physically plausible motion patterns without sacrificing the inherent capabilities of generative models. Experimental results demonstrate that our framework outperforms both traditional codecs and state-of-the-art end-to-end video compression methods under ULB conditions. Furthermore, additional experiments confirm that our approach achieves more precise motion control than existing text-guided methods, paving the way for a novel direction of generative video coding guided by geometric motion modeling.
1 Introduction
One of the core challenges in video coding lies in effectively modeling inter-frame dependencies to reduce temporal redundancy and enhance coding efficiency. While traditional video coding standards (e.g., H.266/VVC (Bross et al. 2021) and AVS3 (Zhang et al. 2019)) have achieved remarkable progress, their reliance on handcrafted motion compensation and transform-quantization modules exhibits limitations in modeling non-rigid motions and nonlocal spatio-temporal dependencies, particularly under Ultra-Low Bitrate (ULB) scenarios with constrained bandwidth. Benefiting from end-to-end optimization strategy, recently emerging deep learning-based video coding approaches (Man et al. 2024; Li, Li, and Lu 2023, 2024; Jiang et al. 2024) have demonstrated promising potential to surpass conventional schemes. However, these methods primarily emphasize pixel-level signal fidelity and suffer from semantic information loss in ULB scenarios due to limited temporal context information. Although recent attempts (Yang, Timofte, and Van Gool 2022; Du, Liu, and Ling 2024; Du et al. 2022) have leveraged GAN (Goodfellow et al. 2020) to improve the perceptual quality, they still adopt similar frameworks as (Li, Li, and Lu 2023, 2024; Jiang et al. 2024; Yang et al. 2020) to model temporal contextual information, which constraints their ability to achieve lower bitrates. Consequently, preserving critical semantic information while enhancing perceptual quality has become a pressing challenge for ULB video compression.
Recent breakthroughs in video generation technology have provided new feasibility for video coding under ULB conditions, catalyzing the emergence of the concept of ”generative video coding” (Chen et al. 2024b). Unlike generation tasks in computer vision, generative video coding primarily aims to achieve content-faithful reconstruction under strict bitrate constraints by leveraging the strong priors of generative models as well as compact spatio-temporal guidance. Despite these theoretical advantages, most existing generative video coding schemes are tailored to specific types of videos, such as human face video (Chen et al. 2024a), human body video (Wang et al. 2022, 2023), and small motion video (Yin et al. 2024), due to the limited capabilities of earlier-stage generative models.
The rapid advancement of video diffusion models (VDMs) (Xing et al. 2024b) has significantly intensified research interest in extending generative video coding to natural video content. Specifically, some recent works (Zhang et al. 2024b; Wan, Zheng, and Fan 2025) attempt to model spatio-temporal information solely based on texts and keyframes from the original video. However, although text serves as an effective modality for representing high-level semantics of video, text-guided generative models often encounter difficulties in reconstructing motion with high perceptual quality and may even produce visibly inferior results in certain cases. In parallel, image-to-video or image-text-to-video models (Xing et al. 2024a; Zhang et al. 2024a) have shown the potential to synthesize motion sequences conditioned on the structural and textual information derived from a pair of keyframes. Nevertheless, these methods still struggle to ensure temporal consistency and visual stability when applied to complex real-world scenarios. The generated outputs may exhibit unexpected artifacts, such as unnatural motion patterns or missing dynamic objects, compromising the temporal semantic fidelity with the original video. To allow for more precise generation, an intuitive idea is to take advantage of low-level motion information, as shown in Figure 1. Unfortunately, employing dense motion representation incurs a higher bitrate cost for VDMs to generate high-fidelity videos. It is essential to find an effective way to extract compact motion guidance which contains key temporal semantic information from the original video.
Figure 1: Examples of motion representation for video (event, optical flow, trajectory and text).
Another important aspect concerns the conditional guidance mechanisms employed in the generative model. Most existing controllable video generation methods utilize ’classifier-free guidance’ (Ho and Salimans 2022) to fine-tune pre-trained video diffusion models, such as Stable Video Diffusion (SVD) (Blattmann et al. 2023) and VideoCraft (Chen et al. 2023b, 2024c), or adopt a ControlNet-like adapter (Zhang, Rao, and Agrawala 2023; Niu et al. 2024) to steer the generation process according to user specifications. However, the loss of critical structural and textural details in compressed keyframes significantly impairs the model’s ability to reconstruct motion-aligned videos, particularly under ULB conditions. Moreover, these pre-trained generative models suffer from limited adaptability in the context of video coding tasks, as the compact temporal guidance (e.g., motion trajectory) may deviate substantially from the original conditioning domains. How to effectively leverage compact guidance information to control the generative model in reconstructing high-fidelity and motion-aligned videos also requires further explored.
In response to the critical challenge of generative video coding under ULB conditions, we propose a Trajectory-guided Generative Video Coding framework. To balance the trade-off between temporal semantic information preservation and bitrate saving, a semantic-aware sparse motion sampling pipeline is first proposed to bridge low-level motion tracking with high-level semantic understanding. Specifically, we track the motion of pixels on a pre-defined grid and classify them into distinct motion instances according to motion patterns. To further improve coding efficiency, a sparse sampling of motion instances is performed based on their semantic importance, representing motion as a set of trajectory points. The variations in trajectory points reflect different motion patterns (such as translation, occlusion, deformation, etc.), which preserves critical temporal semantic information of the original video while significantly reducing the bitrate. To guide the generative model in reconstructing videos aligned with target trajectories without compromising its inherent generative capabilities, we propose a training-free guidance approach. Unlike existing methods that constrain feature maps in intermediate network layers, we design a simple yet effective guidance function and directly impose constraints on the latent space to ensure that the trajectories of target points align with real-world motion paths.
To this end, our principal contributions are summarized as follows:
  • We propose a semantic-aware sparse motion sampling pipe-line tailored for generative video coding, which bridges low-level motion tracking with high-level semantic understanding by encoding motion as sparse trajectory points, significantly reducing bitrate while preserving temporal semantics.
  • A training-free latent space guidance mechanism is designed to enforce trajectory alignment via a lightweight guidance function without compromising the inherent capabilities of the generative model.
  • Extensive experiments demonstrate that T-GVC outperforms both traditional codecs and state-of-the-art deep learning-based methods under ULB conditions, establishing a novel paradigm for efficient semantic-aware video coding in resource-constrained scenarios.
2 Related Work
Figure 2: Overview of our T-GVC framework. On the encoder side, each pair of keyframes and corresponding inter-frames are fed into proposed sparse motion sampler to extract motion trajectories. Subsequently, the keyframes and trajectories are encoded into compact bitstreams. On the decoder side, each decoded keyframe pair is encoded into latent features via VAE encoder. These latent features, combined with zero-initialized latent features, form a latent sequence and concatenated with the initial latent noises as input of VDM. The decoded sparse motion trajectories act as guidance conditions during the inference process to correct the motion in latent space. Ultimately, the clean output is decoded by VAE.
2.1 Generative video coding
Most end-to-end video coding frameworks optimized for signal fidelity (Shi and Lu 2024) face critical challenges in maintaining satisfactory perceptual quality under ULB constraints. Recently, generative approaches have been explored to improve perceptual quality in ULB scenarios, giving rise to the concept of generative video coding (Chen et al. 2024b). Early generative video coding methods primarily targeted specific content types, such as facial or human-centric videos. For instanse, Wang et al. (Wang, Mallya, and Liu 2021) first proposed video synthesis based on reference frames and keypoints for facial video compression. Subsequently, Feng et al. (Feng et al. 2021) incorporated adversarial training with additional modalities to improve visual fidelity, and such a strategy was also adopted in (Oquab et al. 2021; Hong et al. 2022). In (Shukor et al. 2022), Shukor et al. projected video frames into the latent space of StyleGAN (Richardson et al. 2021) for efficient compression, while Chen et al. (Chen et al. 2023a) introduced a dynamic reference refresh mechanism to enhance temporal stability. Beyond facial contents, Wang et al. (Wang et al. 2022) decomposed human-centric videos into texture and structure via contrastive learning and human pose keypoints, and future work (Wang et al. 2023) leveraged Principal Component Analysis (PCA) to compress motion features for more compact human video representations.
For natural videos, early approaches (Yang, Timofte, and Van Gool 2022; Du et al. 2022; Du, Liu, and Ling 2024) employed GANs to enhance perceptual quality under low bitrate conditions. More recently, a framework based on VQ-VAE (Qi et al. 2025) further performed transform coding in the latent space to achieve high realism and high fidelity at ULB. However, constrained by the inherent architectures of conventional frameworks (Li, Li, and Lu 2023, 2024; Jiang et al. 2024; Yang et al. 2020), these methods are difficult to generalize to even lower bitrate scenarios. Inspired by Cross-Modal Compression (CMC) in the field of image compression (Li et al. 2021; Zhang et al. 2023; Gao et al. 2023, 2024), Zhang et al. (Zhang et al. 2024b) designed Cross-Modal Video Compression (CMVC) to explore the potential of applying multimodal large language models for ULB video compression. In CMVC, a video is segmented into multiple semantically coherent clips, each of which is decomposed into content (a keyframe) and motion (text descriptions). On the decoder side, keyframes are reconstructed first, followed by the generation of intermediate frames guided by the extracted ’motion’ information. Similarly, Wan et al. (Wan, Zheng, and Fan 2025) utilized a text-guided video diffusion model for video reconstruction, incorporating a more elaborate keyframe selection strategy. While these approaches are able to preserve essential semantic contents at ULB, they still suffer from issues with temporal consistency when handling complex motion patterns. Overall, despite significant progress in generative video coding for specific content types, extending these techniques to natural video, particularly in dynamic and motion-rich scenes, remains an open and challenging topic.
2.2 Motion-conditioned diffusion model
Text-conditioned video diffusion models (VDMs) have demonstrated remarkable capabilities in synthesizing visually striking content. However, they still struggle to achieve precise outcome control, frequently generating unrealistic sequences that violate physical plausibility. To address this limitation, recent studies have increasingly focused on incorporating explicit motion guidance into diffusion models to improve temporal coherence and realism of generated videos. As one of the pioneering works, MCDiff (Chen et al. 2023c) performed video synthesis conditioned on motion signals by providing the first video frame along with a sequence of stroke motions. A flow completion model was employed to convert sparse inputs into dense motion maps, followed by an auto-regressive strategy to predict subsequent frames. Building upon this idea, DragNUWA (Yin et al. 2023) improved the motion-guided scheme by integrating text, image, and trajectory information into VDMs, enabling more fine-grained control over video content generation. More recently, MotionCtrl (Wang et al. 2024) proposed multiple dedicated modules to manage camera motion and object motion independently, while Tora (Zhang et al. 2024c) first integrated trajectories into Diffusion Transformer (DiT) framework (Peebles and Xie 2023), producing longer videos with better physical realism.
All of the aforementioned guidance approaches can be seamlessly compatible with existing diffusion-based video generation frameworks, enhancing the overall motion modeling capabilities. However, they typically require training new diffusion models or motion control modules from scratch. To mitigate the trade-off between quality and computation cost, more recent studies (Xiao et al. 2024; Zhang et al. 2025) have explored the use of classifier guidance (Dhariwal and Nichol 2021) to control the video generation process in a more efficient manner. Leveraging a PCA-based analysis of VDMs, Xiao et al. (Xiao et al. 2024) revealed the presence of intrinsic motion-aware features within the latent representations of VDMs, and proposed a training-free framework for motion control. Similarly, Zhang et al. (Zhang et al. 2025) extracted motion patterns as soft trajectories from the feature maps of the temporal attention modules in VDMs, and performed classifier guidance (Dhariwal and Nichol 2021) to align the denoising process with the target motion patterns. These approaches have shown the flexibility and effectiveness of the training-free video motion control framework, which provided valuable inspiration for our work.
3 METHODOLOGY
3.1 Overview
The overall structure of the proposed T-GVC framework is shown in Figure 2, which follows the common paradigm of generative video coding. Within the framework, the semantic-aware sparse motion sampling pipeline serves as a sparse motion sampler to capture motion from inter-frames, while the latent space guidance mechanism plays a role in motion correction during the DDIM (Song, Meng, and Ermon 2020) denoising process. In the remainder of this section, we will describe the entire coding pipeline of T-GVC and elaborate on each module. For ease of understanding, we list the main symbols used in this paper in Table 1.
3.2 Keyframe selection and compression
Following (Zhang et al. 2024b), we first segment the source video into multiple clips and select the frame with the smallest pairwise semantic similarity as the keyframe, which is measured via a CLIP-based criteria (Radford et al. 2021). To prevent visual artifacts caused by abrupt scene transitions, we additionally introduce a semantic-aware scene cut detection mechanism based on inter-frame similarity analysis. Specifically, when the semantic similarity score between two consecutive frames falls below a pre-defined threshold, both frames are designated as keyframes to ensure seamless transitions. Therefore, more than one keyframes may be selected within each video clip.
Each pair of consecutive keyframes $\{K_{i},K_{i+1}\}$, together with their associated intermediate frames, constitutes a new clip $\{F_{0},F_{1},...,F_{L-1}\}$. The clip structure serves as the fundamental processing unit for subsequent motion sampling and semantic-consistent video reconstruction. Notably, since keyframes carry critical semantic information, their reconstruction on the decoder side must maintain high perceptual quality even under stringent bitrate constraints, thereby preserving essential textural details and structural features to facilitate subsequent intermediate frame reconstruction. As this work primarily focuses on inter coding, we adopt the off-the-shelf MS-ILLM (Muckley et al. 2023) for keyframe compression, owing to its demonstrated ability to preserve superior perceptual fidelity under aggressive bitrate constraints.
Symbol Description
$F_{i}$ The $i$-th frame in the video clip.
$\mathcal{T}$ Set of motion trajectories.
$\mathcal{T}^{i}$ The $i$-th motion instance (cluster).
$T_{i}\in\mathcal{T}$ The $i$-th trajectory in the trajectory set.
$(x_{j}^{i},y_{j}^{i})$ Spatial coordinates of $T_{i}$ in $F_{j}$.
$V_{j}^{i}\in\{0,1\}$ Visibility of $T_{i}$ at $F_{j}$.
$M_{i}$ Motion mask of $\mathcal{T}^{i}$.
$z_{t}^{i}$ The $i$-th latent feature at time step $t$.
$\tau_{i}^{j}$ Sparse trajectory points of $\mathcal{T}^{j}$ in $F_{i}$.
$L$ Number of frames in the video clip.
Table 1: Summary of symbols used in T-GVC.
3.3 Semantic-aware sparse motion sampling
To bridge low-level motion tracking with high-level semantic understanding, we propose a semantic-aware sparse motion sampling framework that quantifies the semantic importance of different motion trajectories in the original video. The pipeline comprises three stages:
Dense Trajectory Extraction.
Given two keyframes $\{K_{i},K_{i+1}\}$ and their intermediate frames, we design a bidirectional tracking scheme to extract dense motion trajectories with improved temporal consistency. For each clip $\{F_{0},F_{1},...,F_{L-1}\}$ with temporal length $L$, we initialize isometric grids (grid size $=N_{grid}\times N_{grid}$) in the first frame and employ Co-Tracker (Karaev et al. 2024) to trace pixel-wise displacements originating from the grid vertices. The generated forward trajectory sequences over $L$ frames are represented as temporal chains of 3D coordinate points:
$\mathcal{T}_{fwd}=\left\{\left.T_{i}^{fwd}\right|T_{i}^{fwd}={(x_{t}^{i},\ y_{t}^{i},V_{t}^{i})}_{t=0}^{L-1}\right\}$ (1)
To mitigate trajectory loss caused by occlusions or rotation, we reverse the temporal order of each clip and extract the corresponding backward trajectory sequences using the same procedure:
$\mathcal{T}_{bwd}=\left\{\left.T_{j}^{bwd}\right|T_{j}^{bwd}={(x_{t}^{j},y_{t}^{j},V_{t}^{j})}_{t=0}^{L-1}\right\}$ (2)
Subsequently, the backward trajectories are temporally realigned and fused with their forward counterparts to synthesize the final dense trajectories, which holistically preserve the motion characteristics between keyframes:
$\mathcal{T}_{dense}={\mathcal{T}}_{fwd}\cup Flip(\ \mathcal{T}_{bwd}\ )$ (3)
$Flip\left(\cdot\right)$ denotes the reverse operation in the time dimension.
Motion instantiation.
To differentiate between distinct motion patterns, the raw trajectories are clustered using HDBSCAN (McInnes et al. 2017), a hierarchical density-based algorithm particularly effective in handling variable cluster densities and suppressing noise, which is critical for processing real-world videos with irregular motions. In this stage, $T_{i}$ is first transformed into 2D trajectories by excluding points where $V_{t}^{i\ }=0$. Then, we formulate each $T_{i}$ as a spatio-temporal feature vector $f^{i}_{traj}$ as follows:
$f^{i}_{traj}=\left[x^{i}_{0},\ y^{i}_{0},\ \Delta x^{i},\ \Delta y^{i},\ d^{i},\ \bar{\Delta\theta^{i}}\right]$ (4)
[…]
By clustering these features, the corresponding trajectories are categorized into different motion instances $\mathcal{T}^{0},\mathcal{T}^{1},...,\mathcal{T}^{n-1}$ and form the corresponding motion masks, as illustrated in Figure 2.
Semantic importance evaluation.
The extracted dense motion trajectories inherently encapsulate the spatio-temporal dynamics of the source video, enabling the decoder to reconstruct inter-frame semantic relationships through diffusion-based generation (Section 3.4). However, naively encoding these raw trajectories introduces significant bitrate overhead. To reduce inter-frame bitrate while preserving semantically critical motion information, we propose a trajectory sampling strategy driven by semantic loss awareness.
For each motion instance $\mathcal{T}^{i}$, we calculate the semantic importance score of motion $S^{i}_{inter}$ by multiplying the intra-semantic importance score $S^{i}_{intra}$ with the length of the trajectory. The intra-semantic importance score is defined as:
$S^{i}_{\text{intra}}$ $=\sum_{j=0}^{L-2}\big{\|}D(F_{j},F_{j+1})$
$-D\left[P(F_{j},M^{i}),P(F_{j+1},M^{i})\right]\big{\|}_{1}$ (5)
where $D(\cdot)$ denotes the CLIP-based similarity, and $P(F_{j},M^{i})$ represents that the region covered by the motion mask $M^{i}$ in $F_{j}$ is replaced by the surrounding pixels to avoid feature mismatch for CLIP.
We select the motion whose inter-frame semantic importance exceeds the threshold and extract sparse trajectories $\tau^{i}$ through k-means clustering, where the keypoint quantity $K$ for each instance is determined by the semantic importance score $S^{i}_{inter}$ and the quantity of trajectories $N^{i}$:
$K^{i}=min\{(\alpha\cdot S^{i}_{inter}+\beta\cdot\frac{N^{i}}{N_{\text{total}}})\cdot{K_{max}},K_{max}\}$ (6)
$N_{total}$ represents the grid quantity while hyperparameters $\alpha$ and $\beta$ control the weighting ratios of different metrics.
Considering that the video resolution is downsampled by a factor of 8 after projected into the diffusion latent space on the decoder side, the trajectory coordinates are quantized into the same numerical range to ensure alignment with the compressed latent representation. Finally, the initial trajectory data from the first frame, along with the coordinate displacements between adjacent frames are losslessly encoded into a compact bitstream, which would be decoded to guide the reconstruction of the inter-frames on the decoder side.
(a) Reconstructed frame 1
(b) Latent feature map of frame 1 (T=9)
(c) Reconstructed frame 2
(d) Heatmap on latent feature map of frame 2 (T=9)
Figure 3: Illustration of the semantic correlation of latent features (upscaled to the same resolution as the original frame) in the same trajectory across two reconstructed frames. Given the red trajectory point in (b), we plot (d) according to the similarity between the feature on the point and the latent features of frame 2.
3.4 Video reconstruction
Trajectory-based motion guidance.
Inspired by (Xiao et al. 2024; Zhang et al. 2025), we utilize motion trajectories to guide the generation process in a training-free manner by imposing trajectory-aware constraints during the denoising diffusion process. In (Xiao et al. 2024; Zhang et al. 2025), the trajectories are extracted from the feature maps of U-Net blocks in VDM. However, these approaches incur substantial computational overhead during inference, particularly for generative models with a large number of parameters. Moreover, since the motion of generated results is globally controlled, it is difficult to guide the model’s output with fine-grained precision. In contrast, T-GVC directly imposes constraint on the latent noise, leveraging the insight that the latent space in video diffusion contains crucial structural information and semantic correlation. As shown in Figure 3, the region with the highest similarity in Figure 3(a) corresponds to the position of the red point in Figure 3(c), which belongs to the same trajectory as the red point in Figure 3(a) or Figure 3(b).
For each clip of the target video with length $L$, the pipeline on the decoder side begins with decoding compressed keyframes and associated motion trajectories. After that, the keyframes $\{{\hat{K}}_{i},{\hat{K}}_{i+1}\}$ are projected into latent space via VAE (Kingma, Welling et al. 2013) encoder and concatenated with the initial noisy latent $z_{T}=\{z_{T}^{0},\ z_{T}^{1},\ \ldots,\ z_{T}^{L-1}\}$ as conditions, providing the key textual and structural features of the original video.
Let $F\left(z_{t}^{i},\tau_{i}^{j}\right)$ denote the feature of the $i-th$ latent $z_{t}^{i}$ at time step t within the region covered by the sparse trajectories $\tau_{i}^{j}$ of the motion instance $\mathcal{T}^{j}$. To reconstruct the temporal semantic information in the guidance of trajectories, we define a loss function to measure how well the trajectories of $z_{t}$ align with the trajectory guidance, thereby constraining the denoising process:
$\mathcal{L}_{m}=\ \sum_{i=1}^{L-2}\sum_{j=0}^{n-1}\left[\alpha_{i}\cdot S_{0}\left(z_{t}^{i},\tau_{i}^{j}\right)+\beta_{i}\cdot S_{L-1}\left(z_{t}^{i},\tau_{i}^{j}\right)\right]$ (7)
where $S_{k}(z_{t}^{i},\tau_{i}^{j})=\|{F(z_{t}^{k},\tau_{k}^{j})-F(z_{t}^{i},\tau_{i}^{j})}\|_{1}$, captures the non-local similarity of the latent sequence, while $\alpha_{i}$ and $\beta_{i}$ control the level of the similarity.
We could update the output $\epsilon_{\theta}$ of the model in each time step:
${\hat{\epsilon}}_{\theta}{\left(z_{t},t\right)}=\epsilon_{\theta}{\left(z_{t},t\right)}+{s\left(t\right)}\cdot\nabla_{z_{t}}\mathcal{L}_{m}\left({f(z_{t})}\right)$ (8)
where $f(\cdot)$ is the map from noise latent $z_{t}$ to the clean latent $z_{0}$.
Figure 4: The generation process of our framework when the target video clip length exceeds 16 ($\mathbf{\boldsymbol{{L}={21}}}$).
Video generation with variable length.
Most existing VDMs are trained to generate videos of fixed length (e.g., 16 frames). In contrast, our framework accommodates variable intervals between adjacent keyframes. To this end, we further design a variable-length generation scheme to enhance the flexibility of pre-trained VDMs.
Specifically, for each video clip with length $L\leq 16$, we first interpolate the sparse trajectories $\tau$ to $\hat{\tau}$ with length $\hat{L}=16$ and assign corresponding positional markers to the original trajectories. The denoised latent frames at these marked positions are then selected and concatenated to form the final video. For each video clip with length $L>16$, we perform a dual-stage generation strategy where the trajectories are divided into two segments, with each segment guiding one generation stage separately. Figure 4 demonstrates the operational scenario of our framework in processing a video clip beyond 16 frames, with a theoretical maximum capacity of 30 frames. Notably, increasing the number of processing stages facilitates extended video generation with enhanced temporal coherence, while linearly escalating computational demands.
Figure 5: The R-D performance comparison results for HEVC Class B, Class C, UVG and MCL-JCV datasets.
4 Experiments
4.1 Experimental setup
Dataset and metrics.
Following (Wan, Zheng, and Fan 2025), we select HEVC Class B, C (Bossen 2010) as well as UVG (Mercat, Viitanen, and Vanne 2020) and MCL-JCV (Wang et al. 2016) as the test dataset to evaluate the rate-distortion (R-D) performance. All test videos are resized to $512\times 320$ and consist of 96 frames with the original frame rate. For the calculation of rate-distortion metric, the compression rate is quantified in bits per pixel (bpp), while the distortion is measured using LPIPS (Zhang et al. 2018) for perceptual quality and CLIP-SIM (Radford et al. 2021) for semantic similarity. Lower LPIPS and higher CLIP-SIM at the same bitrate represent better coding performance.
Implementation details.
We select the pre-trained DynamiCrafter (Xing et al. 2024a) video diffusion model as our inter-frame decoder. Alternative video generation models designed for open-domain image interpolation or morphing are also compatible with our framework. We remove text prompts from the original model and solely utilize trajectories and keyframes to guide the generation of inter-frames. The number of DDIM denoising steps for each video and the classifier guide scale $s\left(t\right)$ are set to 10 and $30\cdot\sqrt{1-\alpha_{t}}$, respectively. For keyframe compression, we adopt MS-ILLM (Muckley et al. 2023) under ’quality 1’ and ’quality 2’ settings. In addition, given that the video resolution is downsampled to $64\times 40$ by a factor of 8 after projected into latent space, the grid size for sparse motion sampling is configured as 64 to preserve structural consistency. All experiments are conducted on a single NVIDIA RTX 3090 GPU.
Figure 6: Visual quality comparison: ground truth, DCVC-FM, VTM and proposed T-GVC (top to bottom). The reconstructed frames of our framework demonstrates higher perceptual quality at similar bitrates.
4.2 Quantitative and qualitative results
We compare T-GVC with traditional codecs (H.265/HEVC (Sullivan et al. 2012) and H.266/VVC (Bross et al. 2021; Man et al. 2025)) and a neural video codec (DCVC-FM (Li, Li, and Lu 2024)) to evaluate its coding performance. The comparison results with other generative video coding schemes are provided and discussed in Section 4.3. For H.265, we use HM-18.0 with QP = 51, 45 and 39. For H.266, we use VTM-23.7 with QP = 63, 57, 51, 45 and 39. For DCVC-FM, we set the $q$ indexes to 0, 8, 16, 24 and 32, which control the compression level. All codecs above are tested under Low Delay P configuration with intra-period$=$-1.
The R-D curves for the test video classes are presented in Figure 5, where the bitrate of our T-GVC is controlled by adjusting the quality of keyframes. As observed, T-GVC outperforms both traditional codecs and the neural video codec in terms of perceptual reconstruction quality and semantic fidelity at ULB conditions. Figure 6 shows the visual results for the sequence ’$BasketballDrill\_832x480\_50$’ from the HEVC Class C dataset. To ensure a fair comparison, reconstructed sequences from different methods are selected at comparable bitrates. It is observed that the reconstructed videos generated by T-GVC demonstrate superior fidelity in the background regions and key foreground objects (e.g., basketball and the players in red). In contrast, VTM exhibits noticeable blocking artifacts while DCVC-FM suffers from significant high-frequency detail loss, leading to degraded texture quality. It is worth noting that although some texture and structural details in the reconstructed frames might be partially missing or inaccurate to some extent (e.g., the direction of the ball and the appearance of players in green), the overall visual quality remains satisfactory for the ULB scenarios.
Models Settings BD-rate (%)
Text W/o motion bitrate -4.45
W/ motion bitrate 31.92
Trajectory W/o motion bitrate -16.84
W/ motion bitrate -3.06
Text+Trajectory W/o motion bitrate -13.33
W/ motion bitrate 34.97
Table 2: Ablation study on the guidance mechanism.
4.3 Ablation study
To further verify the rationality of our design, we conduct ablation studies on two core components of the T-GVC framework: the guidance mechanism and the motion sampler module, as well as the keypoint quantity within each trajectory instance. Video sequences characterized by large motions, including $ParkScene$, $BasketballDrill$, $PartyScene$, $videoSRC02$ and $videoSRC05$ from HEVC Class B, Class C and MCL-JCV, are selected for evaluation. The coding efficiency is assessed using the Bjøntegaard Delta rate (BD-rate) (Bjontegaard 2001), where PSNR is replaced by $1-LPIPS$. Negative BD-rate values indicate bitrate saving for equivalent perceptual quality. We apply DynamiCrafter with empty prompts (i.e., only guided by keyframes) as the comparison anchor in our experiments. The reported BD-rate results are calculated by averaging all test sequences.
Models Settings BD-rate (%)
Dense W/o motion bitrate -17.88
W/ motion bitrate 188.77
Random W/o motion bitrate -7.95
W/ motion bitrate 15.05
Sparse W/o motion bitrate -16.84
W/ motion bitrate -3.06
Table 3: Ablation study on the sparse motion sampler.
Models Settings BD-rate (%)
Sparse $K_{max}$ = 5 W/o motion bitrate -5.94
W/ motion bitrate 4.12
$K_{max}$ = 10 W/o motion bitrate -13.84
W/ motion bitrate -1.86
$\mathbf{\boldsymbol{K_{max}}}$ = 15 W/o motion bitrate -16.84
W/ motion bitrate -3.06
$K_{max}$ = 20 W/o motion bitrate -10.58
W/ motion bitrate 5.82
Table 4: Ablation study on the keypoint quantity.
Ablation study on the guidance mechanism
In prior generative video coding methods (Zhang et al. 2024b; Wan, Zheng, and Fan 2025), text is commonly adopted as motion guidance. To demonstrate the superiority of our guidance mechanism, we implement three variants of generative frameworks based on DynamiCrafter: one guided by text (denoted as Text), one guided by trajectories (denoted as Trajectory) and the other by both text and trajectories (denoted as Text+Trajectory). Trajectories are sampled through our proposed sparse motion sampler, while text descriptions are generated from the source video using Hunyuan-Large (Sun et al. 2024) without limitation on length. The comparison results are presented in Table 2, where ”W/o motion bitrate” refer to excluding the bitrate consumed by motion-related information when calculating BPP, while ”W/ motion bitrate” indicate that this bitrate is included. It is observed that the trajectory-guided approach consistently achieves better coding performance under both evaluation conditions. In particular, the perceptual quality improvement brought by trajectory guidance makes up for the additional bitrate overhead, offering a distinct advantage over text-based guidance. When overhead is not considered, the trajectory-guided method still outperforms the text-guided method. Combining text and trajectory guidance may enhance texture on simple-motion sequences. Some visual examples in Figure 7 further demonstrate that trajectory guidance can perform more precise motion control, especially in dynamic areas. Text guidance can improve perceptual quality to some extent and combining text and trajectory guidance may enhance texture on simple-motion sequences. However, in more complex cases, text prompts may introduce artifacts that conflict with trajectory-aligned motion, leading to degraded results. For instance, as shown in the second column of Figure 7, the basketball appears behind the net, which violates the physical plausibility.
Figure 7: Visual quality comparison for ablation study on the guidance mechanism.
Ablation study on the motion sampler module
To validate the effectiveness of the semantic-aware motion sampling pipeline proposed in T-GVC, we further conduct an ablation study on the motion sampler module. The ablation results are presented in Table 3. Specifically, ”Sparse” refers to motion trajectories extracted with our semantic-aware motion sampling pipeline. ”Dense” denotes applying dense trajectories as guidance without sampling. ”Random” corresponds to randomly selecting sparse trajectories with $k$ keypoints ($K<K_{max}$) for each motion instance. Although dense trajectories can provide more precise motion representation for video reconstruction, such excessive bitrate overhead is intolerable for the video coding task. In contrast, our motion sampling pipeline achieves significant bitrate reduction in inter-frame motion modeling, while preserving perceptual quality comparable to that of dense trajectory guidance. Compared with a random sampling strategy, our approach thoroughly considers the impact of each motion instance on the video semantics, effectively reducing the bitrate while retaining key motion information.
Ablation study on the keypoint quantity
In the design of T-GVC, $K_{max}$ is a critical term that controls the maximum number of clustering central points of each motion instance. To evaluate its impact on coding performance, we perform ablation studies on $K_{max}$ by setting its value to 5, 10, 15, and 20, respectively. The comparison results are shown in Table 4. As observed, selecting an appropriate number of keypoints is essential for maintaining high-fidelity video reconstruction quality. Theoretically, under unconstrained bitrate conditions, increasing the keypoint quantity would lead to improved reconstruction quality. However, due to the limitations of current pixel-tracking models in terms of accuracy and stability, our method demonstrates a paradoxical phenomenon: excessive keypoints may introduce feature mismatch cascades that impair the inference process of the generative model, ultimately degrading reconstruction fidelity.
5 Conclution and Discussion
In this paper, we propose T-GVC, a novel Trajectory-guided Generative Video Coding framework designed for ULB scenarios. Unlike existing methods that rely heavily on domain-specific priors or high-level text guidance, T-GVC introduces a semantic-aware sparse motion sampling strategy, extracting motion trajectories based on semantic importance to retain key temporal information with minimal bitrate overhead. To further enhance motion realism, a training-free latent guidance mechanism based on trajectory-aligned loss is integrated into the diffusion process, enabling accurate and physically plausible reconstructions. Experimental results demonstrate that T-GVC significantly outperforms traditional codecs and state-of-the-art generative compression methods in terms of both reconstruction quality and motion control precision.